# A Spatiotemporal, Quasi-experimental Causal Inference Approach to Characterize the Effects of Global Plastic Waste Export and Burning on Air Quality Using Remotely Sensed Data

### v2

@@ 3.2 Motivations for Efficient Influence Function-Based Estimation
where $f_{\phi}(D)$ is a flexible (nonparametric) function, with parameters $\phi$. There are several drawbacks of this approach. To begin, even if we allow a flexible relationship between $D$ and $Y$, we are not allowing for flexible relationships between $\mathbf{X}$ and $Y$, nor interactions between the components of $\mathbf{X}$. Importantly, we are not addressing confounding between the dose $D$ and trends in the outcome – for instance, if the overall (or seasonal) trend in air pollution is affected by population density, which is static in our dataset and which we know to be correlated with $D$ (see Section 4.2). To address this, we could make a much more complicated linear regression (e.g., by using splines and more interaction terms), or use a nonparametric machine learning algorithm to model $Y$ given $D,\mathbf{X},B$. However, this outcome model might still be misspecified. This motivates the search for a doubly-robust estimator, which incorporates both a model for the outcome and a model for the propensity score (the probability of receiving treatment), which explicitly addresses confounding, and will produce results that are consistent (reliable in a probabilistic sense) if at least one of these models is correctly specified [^3] While some papers have questioned the likelihood of either the [generalized] propensity score or outcome model being correctly specified in observational studies, the most well-established alternative to doubly-robust methods for continuous exposures, which utilizes matching (Wu et al., 2024), has been shown to perform poorly when the sample size is small (Cork et al., 2025). And in our application, we only observe 356 unique values of the exposure, one per dump site..
We turn to estimators based on the the Efficient Influence Function (EIF). Without going into the technical details (interested readers should see Kennedy (2022) and Schuler and van der Laan (2023)), the EIF for a specific estimand can be used to construct an estimator that is doubly robust and attains the nonparametric efficiency bound (a nonparametric analog of the Cramer-Rao bound), at least asymptotically. In practice, one has to derive the EIF only for a novel estimand, and in many cases this can build closely on past work. Kennedy et al. (2017) provided a doubly-robust estimation method for causal exposure-response curves in observational data settings, which Hettinger et al. (2025) then extended to a DiD framework, including conditions for “multiple robustness” which will be described in the following section. We further extend this multiply-robust estimator to a setting without control locations, which as described in Section 1 is another common type of quasi-experiment. This work contributes to the nascent application of EIF-based estimation techniques in quasi-experimental settings.
@@ 3.3 Estimation
The major difference between our estimator and that of Hettinger et al. (2025) is that instead of modeling the pre-post difference in outcomes, we are just modeling the outcome $Y$ because we are using pre-intervention years rather than separate control locations. This does not substantively change the steps of the proofs in Hettinger et al. (2025) to (1) derive the EIF for the ADT [^4] Technically, they derive the EIF for the integral of $ADT(\delta)$ to satisfy the pathwise differentiability requirement. and (2) show that the resulting estimator is multiply robust. At a high level, after obtaining the EIF, we construct the estimator via the method of estimating equations, which involves setting the EIF equal to zero and solving for $\Psi$ (Schuler and van der Laan, 2023). Given our goal of estimating $\Psi(\delta)$, the relevant parts of this estimator are:
@@
[…]
@@
The estimation of the nuisance functions can be carried out using any flexible modeling technique, e.g., tree ensembles or SuperLearner. Local linear kernel regression (LLKR) is employed for the smooth regression in step 3. The methods and implementation details used in our case study analysis are described in Section 3.5.
@@ 3.4 Uncertainty Quantification
For uncertainty quantification, Hettinger et al. (2025) derive a sandwich variance estimator, but mention various limitations of that approach, including that it relies on a normal approximation, that incorporating uncertainty in nuisance function estimation is challenging, and that extensions would be required to account for correlated data. To avoid these complications, here we use a novel variant of the nonparametric bootstrap to obtain pointwise confidence intervals for the ADT. This approach avoids making asymptotic assumptions in our finite sample setting and is designed to respect the complex spatial and temporal correlation structures in our data. Rather than discrete bootstrapping, where observations are either included in a sample or left out entirely, we use weights drawn from an Exponential distribution to allow for small but nonzero inclusion, which helps to improve the observed support of $D$ in the samples (Hettinger et al., 2025). (Note that these weights are scaled to reflect the observed sample sizes in the pre- and post-intervention time periods.) Sampling one weight per location (dump site) rather than per individual observation accounts for temporal correlation across repeated observations. An important extension we propose, motivated by our case study, is to adapt the weights to account for spatial correlation in addition to temporal.
@@ 3.4.1 Spatial Weighted Bootstrap
    Apply the inverse probability transform to obtain samples with a different (target) marginal distribution, such as the Exponential.
In practice, we can estimate $\Sigma$ by modeling the empirical spatial variogram of the residuals (see the next section for more details on this estimation and explanation of which model’s residuals are used). This approach of generating bootstrap weights from an estimated spatial correlation function is substantiated by Kurisu et al. (2023), who show the mathematical validity of what they term the spatially dependent wild bootstrap (SDWB). While the SDWB paper assumes a Gaussian random field, they note that this assumption can be relaxed at the expense of additional technical complexity. The issue, described in depth by Tsoukalas et al. (2020), is that the magnitude of correlation is not preserved during the mapping from Gaussian random variables to those from another distribution using the inverse probability transform. To end up with the desired spatial correlation in our Exponential weights, we must first estimate an “equivalent correlation” $\Sigma^{*}$ to feed into the Multivariate Normal sampling procedure. Namely, $\Sigma^{*}$ should be specified such that, when the Multivariate Normal samples are inverse probability transformed to generate Exponential weights, the resulting weights have the desired spatial correlation ($\Sigma$). The estimation of $\Sigma^{*}$ can be achieved using numerical integration, which is facilitated by the anySim package in R (Tsoukalas et al., 2020).
@@
    Use $w_{i}$ as weights for each observation from location $i$ in the dataset ($t=1,...,96$), and carry these weights through nuisance function estimation and regression on $D$ to obtain an estimate of the ADT.
@@ 3.4.2 Estimating the Residual Spatial Correlation Function
For convenience, let $\hat{\xi}_{it}$ denote $\hat{\xi}(D_{it},X_{it},B_{it},Y_{it})$. Here, we focus on spatial correlation in the residuals $\hat{\xi}_{it}-\hat{\Psi}_{D}(D_{it})$ for all observations in the treated period $B=1$. (Note that the counterfactual outcome $\hat{\Psi}_{0}$ is constant over space.) Intuitively, we address residual spatial correlation from the final stage of modeling because that is the stage in which our effect of interest is estimated, and it represents the “last opportunity” to explain spatial correlation in the data via modeling. From a more technical perspective, Lee et al. (2017) explain that for doubly-robust methods utilizing LLKR, the model-based uncertainty arising from the second-stage estimation (the LLKR) is greater than that of the first-stage estimation (the nuisance models). Additionally, note that in settings like ours where the exposure of interest is spatially explicit, calculating residuals based on the final effect curve (which has been marginalized over all the other covariates) induces a different spatial structure than if we were to analyze residuals $Y_{it}-\hat{\xi}_{it}$.
@@ 3.4.3 Simulations to Investigate the Spatial Weighted Bootstrap
A pseudo-simulation study was conducted to evaluate the properties of the proposed Spatial Weighted Bootstrap confidence intervals. In Appendix C, we provide full details on the data generating process, the calculation of performance metrics, and the results of this simulation study. Crucially, these simulations illustrate that our Spatial Weighted Bootstrap improves coverage of the confidence intervals relative to a Non-spatial Weighted Bootstrap, across a variety of scenarios.
@@ 3.5 Case Study Data Application
In our case study analysis, we use the Ranger algorithm in the caret package in R to estimate the nuisance functions, using separate models for $\mu_{1},\mu_{0},\pi_{B},\pi_{D}$. Details of these models’ hyperparameter tuning and estimation of $\pi_{D}$ (the GPS) and $f(D|B=1)$ are in Appendix A. However, note that while we referred to a common set of covariates ($\mathbf{X}$) throughout the previous sections, the set of covariates used in the two propensity score models is actually slightly smaller than that described in Section 3.1. The time (month) variable is not included in the model for $\pi_{B}$, because the set of covariates in a flexible machine learning model must not be able to perfectly predict the intervention (Heffernan et al., 2024) – in our case, time steps before and after 1/1/2018. Similarly, the dump site ID is not included in the model for $\pi_{D}$, because the port proximity value is unique to each site. The time variable is also not included in the model for $\pi_{D}$ because our port proximity values are static over time.
@@
To regress $\hat{\xi}$ on $D$ (step 3 in the estimation procedure from Section 3.3), we use the locpol package in R to perform LLKR with a Gaussian kernel. This package utilizes leave-one-out cross-validation (LOO CV) to select the bandwidth parameter for the kernel regression. Therefore, naively applying the method to our dataset, which contains repeated observations (over time) at each unique value of $D$ (one per dump site), selects an inappropriately small bandwidth. The resulting confidence intervals indicate high instability, as shown in Figure S5. To address this issue, we instead apply the LOO CV to each time step separately (across the 356 dump sites), and then calculate the average bandwidth across the time steps – recall that for the LLKR stage, we are just considering the post-treatment time period $B=1$. This average bandwidth value is then used in the LLKR model fit on all time points simultaneously.
@@
where $\Gamma$ is the gamma function, $K_{\nu}$ is the modified Bessel function of the second kind, and $\nu$ is an additional smoothness parameter which is optimized alongside $L$.
The empirical spatial range was $\hat{L}=26.8$km, and the optimal smoothness parameter was $\hat{\nu}=0.35$. To gain intuition for the spatially weighted bootstrap, we also ran an analysis ignoring the possibility of spatial correlation and another using $L=10$km (with the same smoothness $\nu=0.35$), to span the closely-clustered and numerous-fire sites near Jakarta (shown in Figure S4). The upper lefthand plot in Figure S7 illustrates the resulting magnitudes of correlation between sites at different distances, when using 10km and 26.8km as the spatial range parameter. Confidence intervals for the ADT (with or without spatial correlation in the weights) are calculated using 100 bootstrap samples.
@@
For illustrative purposes, we compare the results from our multiply-robust EIF-based method to (i) a plain outcome regression, which is the same as that used inside the multiply-robust estimator, and (ii) a flexible but confounder-naive approach. The latter estimates $E[Y|D=\delta,B=1]-E[Y|B=0]$, where the first expectation uses the same flexible regression on $D$ as the main method (LLKR) and the second expectation is a sample mean. We also re-ran the analysis using 2017 and, separately, 2016 & 2017 as the “treated” period, to check for pre-trends / policy anticipation (whether the effect began before 2018). Conducting such tests is standard practice for quasi-experiments.
Lastly, while dump site size is not directly accounted for in our main analyses, we conducted a sensitivity analysis to assess how this factor might influence our results, as we thought it could be an important predictor of air pollution heterogeneity across open dump sites. As mentioned in Section 2, Global Plastic Watch estimates the area (footprint) of each dump site. In reality, the dump sites are continuously fluctuating in size, but the measurements at any given time point are less reliable than their average over time; there are also frequent gaps in the remotely sensed detection time series. Therefore, we took the area at each site to be the average across all measurements.
@@
We did not adjust for GPW site size as a covariate in the main analysis because the area of the dump sites as measured after 1/1/18 is technically a post-treatment variable which can be affected by treatment, referred to as a “descendant” of treatment (note that only 68% of the dump sites have at least one area estimate prior to 2018, so we would have to exclude many sites, potentially biasing our causal effect curve, in order to include only pre-2018 area as a covariate). Descendants of treatment should not be conditioned on in causal inference analyses (Hernán and Robins, 2020). We ran a sensitivity analysis to investigate the impact of including site size (averaged across 2012-2019) as a covariate. For simplicity, this sensitivity analysis uses the same nuisance function hyperparameters and spatial range (for the bootstrap weights) as the main analysis.

### v3

@@ 3.2 Motivations for Efficient Influence Function-Based Estimation
where $f_{\phi}(D)$ is a flexible (nonparametric) function, with parameters $\phi$. Note that a simplified version of this where $f_{\phi}(D)=\phi D$ has been extensively analyzed, for instance see Callaway et al. (2024). There are several drawbacks of this approach. To begin, even if we allow a flexible relationship between $D$ and $Y$, we are not allowing for flexible relationships between $\mathbf{X}$ and $Y$, nor interactions between the components of $\mathbf{X}$. Importantly, we are not addressing confounding between the dose $D$ and trends in the outcome – for instance, if the overall (or seasonal) trend in air pollution is affected by population density, which is static in our dataset and which we know to be correlated with $D$ (see Section 4.2). To address this, we could make a much more complicated linear regression (e.g., by using splines and more interaction terms), or use a nonparametric machine learning algorithm to model $Y$ given $D,\mathbf{X},B$. However, this outcome model might still be misspecified. This motivates the search for a doubly-robust estimator, which incorporates both a model for the outcome and a model for the propensity score (the probability of receiving treatment), which explicitly addresses confounding, and will produce results that are consistent (reliable in a probabilistic sense) if at least one of these models is correctly specified.
We turn to estimators based on the the Efficient Influence Function (EIF). Without going into the technical details (interested readers should see Kennedy (2022) and Schuler and van der Laan (2025)), the EIF for a specific estimand can be used to construct an estimator that is doubly robust and attains the nonparametric efficiency bound (a nonparametric analog of the Cramer-Rao bound), at least asymptotically. Note that, as we will see below, the EIF does not always construct a doubly-robust estimator (Schuler and van der Laan, 2025), and the EIF may not exist for a particular estimand. In practice, one has to derive the EIF only for a novel estimand, and in many cases this can build closely on past work. Kennedy et al. (2017) provided a doubly-robust estimation method for causal exposure-response curves in observational data settings, which Hettinger et al. (2025a) then extended to a DiD framework, including conditions for “multiple robustness” which will be described in the following section. We further extend this multiply-robust estimator to a setting without control locations, which as described in Section 1 is another common type of quasi-experiment. This work contributes to the nascent application of EIF-based estimation techniques in quasi-experimental settings.
@@ 3.3 Estimation
The major difference between our estimator and that of Hettinger et al. (2025a) is that instead of modeling the pre-post difference in outcomes, we are just modeling the outcome $Y$ because we are using pre-intervention years rather than separate control locations. This does not substantively change the steps of the proofs in Hettinger et al. (2025a) to (1) derive the EIF and (2) show that the resulting estimator is multiply robust. A technical note is that the derivation of the EIF here refers to the integral of $ADT(\delta)$ to satisfy the pathwise differentiability requirement.
At a high level, after obtaining the EIF, we construct the estimator via the method of estimating equations, which involves setting the EIF equal to zero and solving for $\Psi$ (Schuler and van der Laan, 2025). Given our goal of estimating $\Psi(\delta)$, the relevant parts of this estimator are:
@@
$m(D|B=1)=\int_{\mathbb{X}}\mu_{1}(D,\mathbf{x})dP(\mathbf{x}|B=1)=E[Y|B=1,D]$
$f(D|B=1)=\int_{\mathbb{X}}\pi_{D}(D|B=1,\mathbf{x})dP(\mathbf{x}|B=1)$
@@
The estimation of the nuisance functions can be done with flexible modeling techniques subject to some constraints which prevent overfitting. For multiple robustness, $\mu_{1}$ and $\pi_{D}$ must either be fit using methods from uniformly bounded function classes with finite uniform entropy integrals (referred to as the Donsker class) or use cross-fitting / sample splitting, where the functions are estimated on one part of the data and used to predict in a held-out fold. The methods used to fit $\mu_{0}$ and $\pi_{B}$ just have to converge in probability to the truth (as $n\rightarrow\infty$).
Local linear kernel regression (LLKR) is employed for the smooth regression in step 3. The methods and implementation details used in our case study analysis are described in Section 3.5.
@@ 3.4 Uncertainty Quantification
For uncertainty quantification, Hettinger et al. (2025a) derive a sandwich variance estimator, but mention various limitations of that approach, including that it relies on a normal approximation, that incorporating uncertainty in nuisance function estimation is challenging, and that extensions would be required to account for correlated data. To avoid these complications, they use a nonparametric bootstrap approach to obtain pointwise confidence intervals for the ADT. Rather than discrete bootstrapping, where observations are either included in a sample or left out entirely, they use weights drawn from an Exponential distribution to allow for small but nonzero inclusion, which helps to improve the observed support of the continuous exposure (in our notation, $D$) in the samples (Hettinger et al., 2025a). To account for temporal correlation across repeated observations, they sample one weight per location (as opposed to per individual observation). These weights are scaled to reflect the observed sample sizes in each treatment group.
An important extension we propose, motivated by our case study, is to adapt the weights to account for spatial correlation in addition to temporal. We also modify the weights-scaling step to reflect the observed sample sizes in the pre- and post-intervention time periods, as we do not have separate control locations under a “universal” intervention.
Note that our approach to addressing spatial correlation is distinct from the weighted block bootstrap approach in a working paper by Hettinger et al. (2025b), in which they sample a weight per block/neighborhood defined as all adjacent ZIP codes to a focal ZIP code, and then sum the weights assigned within each ZIP code from overlapping blocks. Described below, our approach has the advantages of (a) empirically estimating the residual spatial correlation and (b) avoiding artifacts/inconsistencies stemming from the use of location-areas with irregular sizes and/or shapes, such as two sites $m$ miles apart being considered spatially correlated in some places but not in other places, and ZIP codes with more adjacencies (shared borders) systematically receiving larger weights.
@@ 3.4.1 Spatial Weighted Bootstrap
    Apply the inverse probability transform [^3] Also known as the inverse transformation method, this is a basic technique to generate a sample from any probability distribution given its cumulative distribution function or CDF $F_{X}(x)$. In brief, if $U\sim\text{Uniform}[0,1]$, then we can generate $X=F_{X}^{-1}(U)$. If we start by sampling $Z$ from a distribution with CDF $F_{Z}(z)$, then we can generate $X=F_{X}^{-1}(F_{Z}(Z))$. to obtain samples with a different (target) marginal distribution, such as the Exponential.
In practice, we can estimate $\Sigma$ by modeling the empirical spatial variogram of the residuals (see the next section for more details on this estimation and explanation of which model’s residuals are used). This approach of generating bootstrap weights from an estimated spatial correlation function is substantiated by Kurisu et al. (2023), who show the mathematical validity of what they term the spatially dependent wild bootstrap (SDWB). While the SDWB paper assumes a Gaussian random field, they note that this assumption can be relaxed at the expense of additional technical complexity. The issue, described in depth by Tsoukalas et al. (2020), is that the magnitude of correlation is not preserved during the mapping from Gaussian random variables to those from another distribution using the inverse probability transform. To end up with the desired spatial correlation in our Exponential weights, we must first estimate an “equivalent correlation” $\Sigma^{*}$ to feed into the Multivariate Normal sampling procedure. Namely, $\Sigma^{*}$ should be specified such that, when the Multivariate Normal samples are inverse probability transformed to generate Exponential weights, the resulting weights have the desired spatial correlation ($\Sigma$). The estimation of $\Sigma^{*}$ can be achieved using numerical integration, which is facilitated by the anySim package (Tsoukalas et al., 2020) in R.
@@
    Use $w_{i}$ as weights for each observation from location $i$ in the dataset (t=1,,96t=1,...,96), and carry these weights through nuisance function estimation and regression on $D$ to obtain an estimate of the ADT.
@@ 3.4.2 Estimating the Residual Spatial Correlation Function
For convenience, let $\hat{\xi}_{it}$ denote $\hat{\xi}(D_{it},X_{it},B_{it},Y_{it})$. Here, we focus on spatial correlation in the residuals $\hat{\xi}_{it}-\hat{\Psi}_{D}(D_{it})$ for all observations in the treated period $B=1$. (Note that the counterfactual outcome $\hat{\Psi}_{0}$ is constant over space.) Intuitively, we address residual spatial correlation from the final stage of modeling because that is the stage in which our effect of interest is estimated, and it represents the “last opportunity” to explain spatial correlation in the data via modeling. Note that in settings like ours where the exposure of interest is spatially explicit, calculating residuals based on the final effect curve (which has been marginalized over all the other covariates) induces a different spatial structure than if we were to analyze residuals $Y_{it}-\hat{\xi}_{it}$. From a more technical perspective, Lee et al. (2017) note that for doubly-robust methods utilizing LLKR, the model-based uncertainty arising from the second-stage estimation (the LLKR) is greater than that of the first-stage estimation (the nuisance models). Having “no estimation effect” from the first stage is closely related to the conditions for double (or multiple) robustness, constraining the nuisance functions as mentioned in Section 3.3.
@@ 3.4.3 Simulations to Investigate the Spatial Weighted Bootstrap
A pseudo-simulation study was conducted to evaluate the properties of the proposed Spatial Weighted Bootstrap confidence intervals. In Appendix C, we provide full details on the data generating process (based closely on our real-world dataset), the calculation of performance metrics, and the results of this simulation study.
These simulations illustrate that the Spatial Weighted Bootstrap has consistently higher coverage than the Non-Spatial Weighted Bootstrap, due in part to generating wider confidence intervals. Averaging across all the values of the exposure (port proximity), the Spatial approach has a coverage of 96.5% and the Non-Spatial approach has a coverage of 91.8%. However, both methods tend to systematically have higher-than-nominal coverage in the middle of the exposure distribution and lower coverage on the tails. Across several simulation scenarios, the lowest coverage observed (at any value of the exposure) for the Spatial Weighted Bootstrap is 87% whereas for the Non-Spatial approach it dips down to 82% (see Figure S15). Re-evaluating the coverage after subtracting out the average bias illustrates that some, but not all, of the under-coverage is attributable to the finite sample bias of the LLKR (see Figure S18).
@@ 3.5 Case Study Data Application
In our case study analysis, we use the Ranger algorithm in the caret package (Kuhn and Max, 2008) in R to estimate the nuisance functions, using separate models for $\mu_{1},\mu_{0},\pi_{B},\pi_{D}$. Ranger is an efficient implementation of random forest, which has been shown to work well in tabular data settings like this (Grinsztajn et al., 2022), and also implicitly performs a type of variable selection (James et al., 2022), which is helpful for identifying which of the fixed effects, at the province or individual site level, is informative (Grinsztajn et al., 2022). However, because this algorithm does not conform to the finite uniform entropy integral condition used to prove multiple robustness, we use cross-fitting with k=10 sample splits for $\mu_{1}$ and $\pi_{D}$. (For $\mu_{0}$ and $\pi_{B}$ we do not do cross-fitting to reduce the computational intensity – as it is not needed for multiple robustness – instead carefully tuning hyperparameters as described below.) Note that Donsker class alternatives like Highly Adaptive Lasso (HAL) tend to be substantially more computationally intensive than Ranger, even with cross-fitting. Another important note is that cross-fitting in a spatiotemporal setting like this one is an active area of research. Splitting samples into folds by unit and by time period both have their disadvantages. Interestingly, in a variety of panel data simulations, Fuhr and Papies (2024) found violations of the cross-fitting independence assumption to be largely inconsequential for the accuracy of the effect estimates.
@@
Note that while we referred to a common set of covariates ($\mathbf{X}$) throughout the previous sections, the set of covariates used in the two propensity score models is actually slightly smaller than that described in Section 3.1. The time (month) variable is not included in the model for $\pi_{B}$, because the set of covariates in a flexible machine learning model must not be able to perfectly predict the intervention (Heffernan et al., 2024) – in our case, time steps before and after 1/1/2018. Similarly, the dump site ID is not included in the model for $\pi_{D}$, because the port proximity value is unique to each site. The time variable is also not included in the model for $\pi_{D}$ because our port proximity values are static over time. Note that because of this, we sample observations by site into the cross-fitting folds (in addition to initial hyperparameter tuning).
We carefully tune each Ranger model according to its use in our estimation procedure. For the regression models (all except for $\pi_{B}$, which is a binary classification), we use min.node.size=5 and select mtry (the number of variables to possibly split at in each node) using 5-fold cross-validation. For $\mu_{1}$, we sample observations into these folds randomly (within $\mathcal{O}_{B=1}$) and set mtry to be the value beyond which the $R^{2}$ and RMSE do not substantially improve (a difference within 0.01 for $R^{2}$ and 0.1 for RMSE). For $\pi_{D}$, we instead sample observations by site into the folds to account for repeated observations at each location. For $\pi_{B}$, we sample observations randomly from $\mathcal{O}$ and limit mtry such that the empirical probabilities do not get too large (>>0.99) or too small (<<0.001). Lastly, for $\mu_{0}$, we set the folds to be 2012-2014 vs 2015-2017 to emulate this model’s extrapolation to the $B=1$ years, and select mtry to minimize the bias rather than maximize predictive accuracy, as recommended in Botosaru et al. (2024). For all the models, we use num.trees=500.
@@
To regress $\hat{\xi}$ on $D$ (step 3 in the estimation procedure from Section 3.3), we use the locpol package (Ojeda Cabrera, 2025) in R to perform LLKR with a Gaussian kernel. This package utilizes leave-one-out cross-validation to select the bandwidth parameter for the kernel regression. Therefore, naively applying the method to our dataset, which contains repeated observations (over time) at each unique value of $D$ (one per dump site), selects an inappropriately small bandwidth. Therefore, we modified this package’s bandwidth selection function to do leave-one-location-out cross-validation where all repeated observations for a given location are assigned to the same fold, as has been done in other kernel regression analyses with longitudinal data (Wang et al., 2024).
@@
where $\Gamma$ is the gamma function, $K_{\nu}$ is the modified Bessel function of the second kind, and $\nu$ is an additional smoothness parameter which is optimized alongside $L$. The empirical spatial range was $\hat{L}=3.85$km, and the optimal smoothness parameter was $\hat{\nu}=5$. Figure 3 illustrates the resulting magnitude of correlation for sites at different distances apart.
Figure 3: The spatial correlation function (Matérn) fit to the residuals from the LLKR stage.
Confidence intervals for the ADT (with or without spatial correlation in the weights) are calculated using 100 bootstrap samples. In our simulation study, increasing the number of bootstrap replicates to 200 did not substantially change the confidence intervals’ coverage, as shown in Figure S17.
@@
For illustrative purposes, we compare the results from our multiply-robust EIF-based method to (i) a plain outcome regression, which is the same as that used inside the multiply-robust estimator, followed by regression on $D$ using LLKR, and (ii) a flexible but confounder-naive approach. The latter estimates $E[Y|D=\delta,B=1]-E[Y|B=0]$, where the first expectation uses LLKR to regress on $D$ and the second expectation is a sample mean. We also re-ran the analysis using 2017 and, separately, 2016 & 2017 as the “treated” period, to check for pre-trends / policy anticipation (whether the effect began before 2018). Conducting such tests is standard practice for quasi-experiments; Heffernan et al. (2024) demonstrated this technique while utilizing past years as controls.
This pre-trends validation approach connects to our stated assumptions in several ways. Clearly, anticipation of a policy being implemented would violate A1 (arrow of time). A5, on the other hand, would be violated if there was some unmeasured time-varying factor causing PM2.5 to be systematically different pre- and post-2018, in the absence of China’s ban. If the pre-trends analysis finds an effect of treatment in 2017 (prior to treatment initiation), this could also indicate that some important time-varying factors influencing PM2.5 are not being captured by our measured confounders, in violation of A5. (A policy implemented in 2017 could fit this criteria, as noted in our Discussion.) However, the absence of an effect in the pre-trends test only offers evidence of plausibility of A5 but does not formally validate it because we cannot test for unmeasured time-varying confounding between $B=0$ and $B=1$, instead relying on domain knowledge. Also, note that because our pre-trends test is estimating the ADT, it is inherently testing conditional ignorability in expectation.
@@
Lastly, we conducted a sensitivity analysis to assess the potential influence of dump site size, as it could be an important predictor of air pollution heterogeneity across sites. As mentioned in Section 2, Global Plastic Watch estimates the area (footprint) of each dump site. Although time-varying estimates are provided, these measurements are less reliable than their average over time and have frequent gaps, so we used each site’s average area across available estimates (2012-2019). We did not adjust for dump site size in the main analysis because area measured after 1/1/18 is a post-treatment variable which can be affected by treatment, referred to as a “descendant” of treatment, which should not be conditioned on in causal inference analyses (Hernán and Robins, 2020). Note that only 68% of the dump sites have at least one area estimate prior to 2018, so we would have to exclude many sites, potentially biasing our causal effect curve, in order to include only pre-2018 area as a covariate.
For simplicity, this sensitivity analysis as well as the aforementioned outcome regression, confounder-naive approach, and pre-trends tests use the same nuisance function hyperparameters and Spatial Weighted Bootstrap with the same Matérn parameters as our main analysis.